OpenStudy (anonymous):

Simplify the rational expressionx^2+2x-8/x^2+5x-14

OpenStudy (highlyeducated87):

x+4/x+7

OpenStudy (highlyeducated87):

need an explanation

OpenStudy (anonymous):

thanks i got it

OpenStudy (highlyeducated87):

x2 + 2x - 8 = ( x - 2 )( x + 4 ) = x + 4 x2 + 5x - 14 ( x - 2 )( x + 7 ) x + 7 The first thing you have to do is factor the numerator and the denominator. Then cancel off any common factors. Then you get your answer.